OpenStudy (anonymous):

1/(x-4) + 1/(x+1) = (x^2-6)/(x^2-3x-4), solve for x.

OpenStudy (anonymous):

notice that x^2-3x-4=(x-4)(x+1) and x can not be 4 or -1 (why?)

OpenStudy (anonymous):

Because it can't equal 0?

OpenStudy (anonymous):

writte exactly the same on www.wolframalpha.com ....

OpenStudy (anonymous):

remember u cannot divide by "0"

OpenStudy (anonymous):

@alienshe because the Denominator of a fraction can not be 0

OpenStudy (anonymous):

now multiply both sides by (x-4)(x+1) and try to solve it sure u can do it

OpenStudy (anonymous):

I'm trying to do it right now and I'm not sure what to do next. 2x-x^2=-9

OpenStudy (anonymous):

^ What I have now.

OpenStudy (anonymous):

u did good there is a little mistake check it again u should get 2x-x^2=-3

OpenStudy (anonymous):

I forgot to change the sign when I moved the -3. I do that too often...

OpenStudy (anonymous):

What do I do now?...

OpenStudy (anonymous):

ok now u have a quadratic equation x^2-2x-3=0 u can solve it by factoring or solving directly with quadratic equation formula

OpenStudy (anonymous):

Why is it -2x and not +2x?

OpenStudy (anonymous):

i moved 2x and -x^2 to the other side

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

Oh. Yes. I moved -3 to the other side. Your way makes more sense, though.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the standard form of a quadratic equation is \[ax^2+bx+c=0\] now u tell me what are a , b and c for our equation?

OpenStudy (anonymous):

a=1, b=-2, c=-3.

OpenStudy (anonymous):

I solved it by factoring and got x=3, -1.

OpenStudy (anonymous):

I know the answer can't be -1. So x=3

OpenStudy (anonymous):

very well

OpenStudy (anonymous):

Thank you for your help! :) I really appreciate it.

OpenStudy (anonymous):

u did it yourself.